OpenStudy (anonymous):

Identify the axis of symmetry for the graph of f(x) = -(x - 4)2 - 6. x=?

OpenStudy (anonymous):

x = 4 ?

OpenStudy (anonymous):

x = -7

OpenStudy (anonymous):

how'd you guys get those answers?

OpenStudy (anonymous):

i may be wrong just rearranged to find x see what others get though

OpenStudy (y2o2):

You get the answer by finding the vertex of the function Vx = -b/2a so by expanding the bracket you get f(x) = -x^2+8 x-22 so Vx = -8 / -2 = 4 so equation of axis of symmetry is x = 4

OpenStudy (anonymous):

the axis of symmetry is the vertex The equation above is in vertex form y = a(x - h) + k where h is the x value of the vertex and k is the y value of the vertex h is always opposite of the sign while k isnt so since your equation is f(x) = -(x - 4)^2 - 6 your vertex is (4,-6). The axis of symmetry is your x value so x = 4

OpenStudy (anonymous):

thanks <33

OpenStudy (anonymous):

anytime <33

OpenStudy (anonymous):

<333

OpenStudy (anonymous):

<3333

OpenStudy (y2o2):

<33333333333333333333333333 :P

OpenStudy (anonymous):

LOL <33333333

OpenStudy (anonymous):

LOL omg <33